# Automorphisms of free metabelian Lie algebras

### v2

@@ 3. Jacobian matrices
be the degree grading of $U$, i.e., $U_{k}$ is the linear span of all homogeneous of degree $k$ monomials in $y_{1},y_{2},\ldots,y_{n}$. If
@@
Let $T_{n}$ be a free right $U$-module with basis $t_{1},\ldots,t_{n}$. Turn the direct sum
@@
is the free metabelian Lie algebra $M_{n}$ with the free generators $x_{1},\ldots,x_{n}$. In the case of groups the corresponding embedding is called the Magnus embedding [23, 27] and in the case of Lie algebras it follows from Shmel’kin’s Theorem on wreath products [56].
@@
The statement of the next lemma is well known (see [62, Lemma 5],[64]).
@@
$(1)$ $a=\partial(f)$ for some $f\in[M_{n},M_{n}]$;
$(2)$ $Ya=y_{1}a_{1}+\ldots+y_{n}a_{n}=0$ where $Y=(y_{1},\ldots,y_{n})$;
$(3)$ $a=\sum_{i<j}(e_{i}y_{j}-e_{j}y_{i})a_{ij},\ a_{ij}\in U$.
@@
Consequently, $(1)$ implies $(3)$. Obviously, $(3)$ implies $(2)$ since $Y(e_{i}y_{j}-e_{j}y_{i})=0$ for all $i<j$.
Suppose that $(2)$ holds. Then $a_{n}=y_{1}b_{1}+\ldots+y_{n-1}b_{n-1}$ for some $b_{j}\in U$. Set $g=[x_{n},x_{1}]b_{1}+\ldots+[x_{n},x_{n-1}]b_{n-1}$. Then $\frac{\partial g}{\partial x_{n}}=a_{n}$ and the last component of $a^{\prime}=a-\partial(g)$ is zero. Moreover, if $a^{\prime}=(a_{1}^{\prime},\ldots,a_{n-1}^{\prime},0)$ then $y_{1}a_{1}^{\prime}+\ldots+y_{n-1}a_{n-1}^{\prime}=0$. Leading an induction on the number of nonzero components of $a$, we may assume that there exists $h\in[M_{n},M_{n}]$ such that $a^{\prime}=\partial(h)$. Then $a=\partial(g+h)$. $\Box$
Every endomorphism $\phi$ of $M_{n}$ induces an endomorphism of $Y_{n}=M_{n}/[M_{n},M_{n}]$, which can be uniquely extended to an endomorphism of $U=U(Y_{n})$. This endomorphism will be denoted again by $\phi$ for simplicity of notation. If $f\in U$, then the image $\phi(f)$ will be denoted by $f^{\phi}$.
@@
(6) $J(\phi\circ\psi)=J(\phi)J(\psi)^{\phi}.$
The statement of the next lemma is also well known (see [58],[62, Theorem 2]).
@@
Here is a description of Chein automorphisms.
Corollary 1.
@@
Proof. Let $\phi=(x_{1}+f,x_{2},\ldots,x_{n})$ where $f\in M_{n}^{2}$. The Jacobian matrix of $\phi$ is $J(\phi)=I+[\partial(f)\ 0\ldots 0]$. This matrix is invertible if and only if $1+\frac{\partial f}{\partial x_{1}}$ is invertible in $U$ or, equivalently, $\frac{\partial f}{\partial x_{1}}=0$. We show that $\frac{\partial f}{\partial x_{1}}=0$ if and only if $f$ belongs to the ideal generated by all $x_{s},x_{t}]$, where $s,t\neq 1$.
The basis (4) of $M_{n}^{2}$ is convenient to work with $\frac{\partial f}{\partial x_{n}}=0$. Any $f\in M_{n}^{2}$ can be written in the form
@@
then the substitution $y_{1}=\ldots=y_{n-2}=0$ gives that $a_{n-1}=0$. Repeated substitution $y_{1}=\ldots=y_{n-3}=0$ gives $a_{n-2}=0$. Continuing the same discussions, we get $a_{n-1}=a_{n-2}=\ldots=a_{1}=0$. Threfore $\frac{\partial f}{\partial x_{n}}=0$ if and only if $f=f^{\prime}$. $\Box$
Consider the set of Chein automorphisms of the form
(7) $(x_{1}+[x_{2},x_{3}]a,x_{2},\ldots,x_{n}),\ a\in U.$
These automorphisms are elementary if $a$ does not depend on $y_{1}$.
In this paper we denote by $\mathrm{Mat}_{n}(U)$ the set of all square matrices of order $n$ over $U$ (since the more standard notation $M_{n}(U)$ can be confused with the notation of the free metabelian Lie algebra $M_{n}$). As usual, denote by $GL_{n}(U)$ and $SL_{n}(U)$ the general linear group and the special linear group of matrices of order $n$ over $U$, respectively. The identity matrix of order $n$ is already denoted by $I$.
In Section 2 we mentioned that the group $GL_{n}(K)$ will be identified with the group of all linear automorphisms of $M_{n}$. We meant that each linear automorphism $\alpha$ will be identified with its Jacobian matrix $J(\alpha)$. This means that $\alpha$ is a matrix and instead of $J(\alpha)$ in our calculations we will write $\alpha$.
@@
$(a)$ If $\phi,\psi\in\mathrm{IAut}(M_{n})$ then
$J(\phi\psi)=J(\phi)J(\psi);$
$(b)$ If $\phi\in\mathrm{IAut}(M_{n})$ and $\alpha\in GL_{n}(K)$ then
$J(\alpha\phi\alpha^{-1})=\alpha J(\phi)^{\alpha}\alpha^{-1};$
$(c)$ The map
$J:\mathrm{IAut}(M_{n})\to SL_{n}(U)$
is an injective homomorphism of groups and $G=J(\mathrm{IAut}(M_{n}))$ is the set of all matrics from $SL_{n}(U)$ of the form
$I+A,$
where $A\in\mathrm{Mat}(U)$ is a matrix such that $YA=0$.
Proof. Since $\phi$ induces an identity automorphism on $U$, $(a)$ is a direct consequence of (6). Using (6) we also get
$J(\alpha\phi\alpha^{-1})=J(\alpha(\phi\alpha^{-1}))=J(\alpha)J(\phi\alpha^{-1})^{\alpha}=J(\alpha)(J(\phi)J(\alpha^{-1})^{\phi})^{\alpha}=J(\alpha)(J(\phi)J(\alpha^{-1}))^{\alpha}$
$=J(\alpha)J(\phi)^{\alpha}J(\alpha^{-1})^{\alpha}=J(\alpha)J(\phi)^{\alpha}J(\alpha^{-1})=J(\alpha)J(\phi)^{\alpha}J(\alpha)^{-1}=\alpha J(\phi)^{\alpha}\alpha^{-1}.$
In this calculation we again used that $\phi$ induces an identity automorphism on $U$, $J(\alpha^{-1})\in GL_{n}(K)$ is a constant matrix, and that $J(\alpha^{-1})=J(\alpha)^{-1}$ by (6).
If $\phi\in\mathrm{IAut}(M_{n})$, then
$\phi=(x_{1}+f_{1},x_{2}+f_{2},\ldots,x_{n}+f_{n}),$
where $f_{i}\in[M_{n},M_{n}]$ for all $i$, and
$J(\phi)=I+A,\ A=\left[\partial(f_{1})\ \ldots\ \partial(f_{n})\right].$
We get $YA=0$ by Lemma 2. If $J(\phi)=I$ then $\partial(f_{i})=0$ for all $i$. Notice that for any $f\in[M_{n},M_{n}]$ we get $f=(t_{1},\ldots,t_{n})\partial(f)$ by the definition of $\partial(f)$. Consequently, $f_{i}=(t_{1},\ldots,t_{n})\partial(f_{i})=0$. This means $\phi=\mathrm{id}$, i.e., $J$ is injective.
Conversely, let $A\in\mathrm{Mat}_{n}(U)$ such that $YA=0$. By Lemma 2 there exist $g_{1},g_{2},\ldots,g_{n}\in[M_{n},M_{n}]$ such that $A=\left[\partial(g_{1})\ \ldots\ \partial(g_{n})\right]$. Set $\psi=(x_{1}+g_{1},x_{2}+g_{2},\ldots,x_{n}+g_{n})$. Then $J(\psi)=I+A$. If $I+A\in SL_{n}(U)$ then $\psi$ is an automorphism by Lemma 3 and $\psi\in\mathrm{IAut}(M_{n})$. $\Box$
4. Matrix notations
Using Lemma 4, we will represent automorphisms from $\mathrm{IAut}(M_{n})$ by their Jacobian matrices and make calculations in the language of Jacobian matrices. Results formulated in the language of matrices will be also reformulated in the language of automorphisms.
Thus $G$ represents the set of all Jacobian matrices of automorphisms from $\mathrm{IAut}(M_{n})$. Set $T=J(\mathrm{TAut}(M_{n})\cap\mathrm{IAut}(M_{n}))$ and $S=J(\mathrm{ATAut}(M_{n})\cap\mathrm{IAut}(M_{n}))$. We have
$T\subseteq S\subseteq G.$
In order to prove that $\mathrm{ATAut}(M_{n})=\mathrm{Aut}(M_{n})$ it is sufficient to prove that $\mathrm{ATAut}(M_{n})\cap\mathrm{IAut}(M_{n})=\mathrm{IAut}(M_{n})$ since both groups include all linear automorphisms. Thus, it is sufficient to prove that $S=G$.
Let $E_{ij}$ a square matrix with $1$ in the $(i,j)$ position and zeros elsewhere. The order of $E_{ij}$ is usually clear from the context. Recall that
$E_{ij}E_{kl}=\delta_{jk}E_{il},$
where $\delta_{jk}$ is the Kronecker delta. A matrix of the form
$I+aE_{ij},\ i\neq j,\ a\in U,$
is called elementary.
If $i<j$, then for any $k$ denote by $F_{ijk}$ the $n\times n$ matrix with $y_{j}$ in position $(i,k)$, $y_{i}$ in position $(j,k)$, and zeros in other places, that is,
$F_{ijk}=y_{j}E_{ik}-y_{i}E_{jk}.$
In other words, the $k$th column of $F_{ijk}$ is equal to $\partial([x_{i},x_{j}])=e_{i}y_{j}-e_{j}y_{i}$, and the rest are zeros. It is easy to check that
$F_{ijk}F_{stm}=\left\{\begin{array}[]{ccc}y_{t}F_{ijm}&{\rm if}&k=s,\\ -y_{s}F_{ijm}&{\rm if}&k=t,\\ 0&{\rm if}&k\neq s,t\\ \end{array}\right$.,
$E_{ij}F_{kst}=y_{s}E_{ij}E_{kt}-y_{k}E_{ij}E_{st}=\left\{\begin{array}[]{ccc}y_{s}E_{it}&{\rm if}&j=k,\\ -y_{k}E_{it}&{\rm if}&j=s,\\ 0&{\rm if}&j\neq k,s\\ \end{array}\right$.,
and
$F_{kst}E_{ij}=y_{s}E_{kt}E_{ij}-y_{k}E_{st}E_{ij}=\left\{\begin{array}[]{ccc}F_{ksj}&{\rm if}&t=i,\\ 0&{\rm if}&t\neq i\\ \end{array}\right.$.
We refer to these three formulas as the multiplication tables but often use them without reference. Properties of the product of matrices of types
$I+A+B=(I+A)(I+B),$
which is true if $AB=0$, and
$I+(a+b)C=(I+aC)(I+bC),$
which is valid if $C^{2}=0$, to simplify terminology, will be called additivity.
For example, if $f=\sum_{i<j}[x_{i},x_{j}]a_{ij}$, then the Jacobian matrix of the endomorphism (5) is $I+\sum_{i<j}a_{ij}F_{ij1}$, and, by additivity,
$I+\sum_{i<j}a_{ij}F_{ij1}=\prod_{i<j}(I+a_{ij}F_{ij1}).$
Moreover,
$I+(a+b)F_{ij1}=(I+aF_{ij1})(I+bF_{ij1}).$
Consider the set of matrices of the form
(11) $I+aF_{231},\ a\in U.$
These matrices are Jacobian matrices of Chein automorphisms (7).
Here are two remarks:
$(a)$ If $A\in G$ and $\alpha\in GL_{n}(K)$ then $\alpha A^{\alpha}\alpha^{-1}\in G$ by Lemma 4(b)(b). This matrix will be called the linear conjugate of $A$ by $\alpha$. In many places, conjugation by linear automorphisms helps to “linearize” automorphisms [73].
$(b)$ If $J(\phi)=I+A\in S$ and $A$ is a homogeneous matrix, then, usually, it is easy to show that $I+\mu A\in S$ for any $\mu\in K$. This can be done by adding some coefficients in the proofs. For example, this can be done when conjugating matrices (11) using $(x_{1},\mu x_{2},x_{3},\ldots,x_{n})$. For this reason we ignore the difference between the coefficients of this type.
Corollary 2.
$(a)$ The group $T$ is generated by all linear conjugates of matrices of the form (11) where $a$ does not depend on $y_{1}$.
$(b)$ The group $S$ is generated by all linear conjugates of matrices of the form (11).
Proof. By Lemma 1, every automorphism $\mathrm{TAut}(M_{n})\cap\mathrm{IAut}(M_{n})$ is a product of linear conjugates of automorphisms of the form (5). Notice that every automorphism of the form (5) is itself a product of linear conjugates of automorphisms of the form (7) with $a\in K[y_{2},\ldots,y_{n}]$. Then $J(\tau)$ is a matrix of the form (11). If $\alpha\in GL_{n}(K)$ then
$J(\alpha\tau\alpha^{-1})=\alpha J(\tau)^{\alpha}\alpha^{-1}$
by Lemma 4, that is, $J(\alpha\tau\alpha^{-1})$ is the linear conjugate of $J(\tau)$ by $\alpha$. If $\phi,\psi\in\mathrm{IAut}(M_{n})$ then $J(\phi\psi)=J(\phi)J(\psi)$ by the same lemma. This proves $(a)$.
The statement $(b)$ can be checked similarly. $\Box$
5. Tame automorphisms
In this section we assume that either $n\geq 4$ and $K$ is a field of characteristic $\neq 3$ or $n\geq 5$.

### v3

@@ 3. Chein automorphisms
be the degree grading of $U$, i.e., $U_{k}$ is the linear span of all homogeneous monomials in $y_{1},y_{2},\ldots,y_{n}$ of degree $k$. If
@@
Let $T_{n}$ be the free right $U$-module with basis $t_{1},\ldots,t_{n}$. Turn the direct sum
@@
is the free metabelian Lie algebra with free generators $x_{1},\ldots,x_{n}$ [19, 23, 51]. We identify this algebra with $M_{n}$.
@@
The statement of the next lemma is well known [56, 58].
@@
$(a)$ $a=\partial(f)$ for some $f\in[M_{n},M_{n}]$;
$(b)$ $Ya=y_{1}a_{1}+\ldots+y_{n}a_{n}=0$ where $Y=(y_{1},\ldots,y_{n})$;
$(c)$ $a=\sum_{i<j}(e_{i}y_{j}-e_{j}y_{i})a_{ij},\ a_{ij}\in U$.
@@
Consequently, $(a)$ implies $(c)$. Obviously, $(c)$ implies $(b)$ since $Y(e_{i}y_{j}-e_{j}y_{i})=0$ for all $i<j$.
Suppose that $(b)$ holds. Let $a_{i}=-y_{n}b_{i}+a_{i}^{0}$, where $a_{i}^{0}\in K[y_{1},\ldots,y_{n-1}]$ for all $1\leq i\leq n-1$. Then $(b)$ implies that $a_{n}=y_{1}b_{1}+\ldots+y_{n-1}b_{n-1}$. Set $g=[x_{n},x_{1}]b_{1}+\ldots+[x_{n},x_{n-1}]b_{n-1}$. We get $\frac{\partial g}{\partial x_{n}}=a_{n}$ and the last component of $a^{\prime}=a-\partial(g)$ is zero. Moreover, if $a^{\prime}=(a_{1}^{\prime},\ldots,a_{n-1}^{\prime},0)$ then $y_{1}a_{1}^{\prime}+\ldots+y_{n-1}a_{n-1}^{\prime}=0$. Leading an induction on the number of nonzero components of $a$, we may assume that there exists $h\in[M_{n},M_{n}]$ such that $a^{\prime}=\partial(h)$. Then $a=\partial(g+h)$. $\Box$
Every endomorphism $\phi$ of $M_{n}$ induces an endomorphism of $Y_{n}=M_{n}/[M_{n},M_{n}]$, which can be uniquely extended to an endomorphism of $U=U(Y_{n})$. This endomorphism will be denoted again by $\phi$ for simplicity of notation. If $f\in U$, then the image $\phi(f)$ will be denoted by $f^{\phi}$ for convenience.
@@
$J(\phi\psi)=J(\phi)J(\psi)^{\phi}.$
Consequently, if $\phi$ is an automorphism, then $J(\phi)$ is invertible. The converse is also true [53, 56].
@@
Proof. Suppose that $J(\phi)$ is invertible. Without loss of generality, we may assume that
$\phi=(x_{1}+f_{1},\ldots,x_{n}+f_{n}),\ \ f_{1},\ldots,f_{n}\in[M_{n},M_{n}].$
Then
$J(\phi)=I+A,\ \ A=[\partial(f_{1}),\ldots,\partial(f_{n})].$
We get
$J(\phi)^{-1}=I+B,B=\sum_{i\geq 1}(-A)^{i}.$
By Lemma 1, we have $YA=0$. Consequently, $YB=0$. Again, by Lemma 1, there exist $g_{1},\ldots,g_{n}\in[M_{n},M_{n}]$ such that
$B=[\partial(g_{1}),\ldots,\partial(g_{n})].$
Let $\psi=(x_{1}+g_{1},\ldots,x_{n}+g_{n})$. Then
$J(\psi)=I+B=J(\phi)^{-1}$
and
$J(\phi\psi)=J(\phi)J(\psi)^{\phi}=J(\phi)J(\psi)=I.$
Consequently, $\phi\psi=\mathrm{Id}$ and $\phi$ is an automorphism. $\Box$
We now give a description of Chein automorphisms of the form (3).
Lemma 3.
@@
Proof. Let $\phi=(x_{1}+f,x_{2},\ldots,x_{n})$ where $f\in[M_{n},M_{n}]$. The Jacobian matrix of $\phi$ is
$J(\phi)=I+[\partial(f)\ 0\ldots 0].$
This matrix is invertible if and only if $1+\frac{\partial f}{\partial x_{1}}$ is invertible in $U$ or, equivalently, $\frac{\partial f}{\partial x_{1}}=0$. We want to show that $\frac{\partial f}{\partial x_{1}}=0$ if and only if $f$ belongs to the ideal generated by all $x_{s},x_{t}]$, where $s,t\neq 1$.
The basis (1) of $M_{n},M_{n}]$ is convenient to work with $\frac{\partial f}{\partial x_{n}}=0$. Any $f\in[M_{n},M_{n}]$ can be written in the form
@@
then the substitution $y_{1}=\ldots=y_{n-2}=0$ gives that $a_{n-1}=0$. Repeated substitution $y_{1}=\ldots=y_{n-3}=0$ gives $a_{n-2}=0$. Continuing the same discussions, we get $a_{n-1}=a_{n-2}=\ldots=a_{1}=0$. Therefore $\frac{\partial f}{\partial x_{n}}=0$ if and only if $f=f^{\prime}$. $\Box$
4. Tame automorphisms
In this section we always assume that either $n\geq 4$ and $K$ is a field of characteristic $\neq 3$ or $n\geq 5$.
Theorem 1.
Every Chein automorphism of $M_{n}$ of lower degree $\geq 4$ is tame.
Proof. It suffices to prove that every Chein automorphism of the form
$C(a)=(x_{1}+[x_{2},x_{3}]a,x_{2},\ldots,x_{n}),\ a\in U,$
with $\mathrm{ldeg}(a)\geq 2$ is tame. Since
$C(a+b)=C(a)C(b),\ \ a,b\in U,$
we may assume that $a=\gamma y_{1}^{i_{1}}y_{2}^{i_{2}}\ldots y_{n}^{i_{n}}$, $0\neq\gamma\in K$. If $i_{1}=0$ then $C(a)$ is elementary. Assume that $i_{1}\geq 1$.
Case 1. Suppose that $i_{j}\geq 1$ for some $2\leq j\leq n-1$. Consider the elementary automorphisms
$\phi=(x_{1},\ldots,x_{n-1},x_{n}+[x_{2},x_{3}]y_{1}^{i_{1}}\ldots y_{j}^{i_{j}-1}),$
$\psi=(x_{1}-\gamma[x_{j},x_{n}]y_{j+1}^{i_{j+1}}\ldots y_{n}^{i_{n}},x_{2},\ldots,x_{n}).$
Direct calculations give that $\phi,\psi]=C(a)$ is tame.
Case 2. Let $i_{2}=i_{3}=\ldots=i_{n-1}=0$ and $i_{n}\neq 0$. Then $a=\gamma y_{1}^{i_{1}}y_{n}^{i_{n}}$. We can make $\gamma=1$ by conjugating $C(a)$ with $(x_{1},\gamma^{-1}x_{2},x_{3}\ldots,x_{n})$. Then assume that $a=y_{1}^{i_{1}}y_{n}^{i_{n}}$. We have
(5) $[x_{2},x_{3}]a=[x_{2},x_{n}]y_{1}^{i_{1}}y_{3}y_{n}^{i_{n}-1}+[x_{n},x_{3}]y_{1}^{i_{1}}y_{2}y_{n}^{i_{n}-1}$
by the Jacobi identity. If $i_{n}\geq 2$ then
$\phi=(x_{1}+[x_{2},x_{n}]y_{1}^{i_{1}}y_{3}y_{n}^{i_{n}-1},x_{2},\ldots,x_{n})=(x_{1}+[x_{2},x_{3}]y_{1}^{i_{1}}y_{3}^{i_{n}-1}y_{n},x_{2},\ldots,x_{n})^{(3n)}$
and
$\psi=(x_{1}+[x_{n},x_{3}]y_{1}^{i_{1}}y_{2}y_{n}^{i_{n}-1},x_{2},\ldots,x_{n})=(x_{1}+[x_{2},x_{3}]y_{1}^{i_{1}}y_{2}^{i_{n}-1}y_{n},x_{2},\ldots,x_{n})^{(2n)}$
are tame by Case 1. Therefore $C(a)=\phi\psi$ is tame.
If $i_{n}=1$, then set $\alpha=(x_{1},x_{2}+x_{n},x_{3},\ldots,x_{n})$. We have
$\tau=(x_{1}+[x_{2},x_{3}]y_{1}^{i_{1}}y_{2},x_{2},\ldots,x_{n})$
is tame by Case 1. Direct calculations give that
$[\alpha,\tau]=(x_{1}+([x_{2},x_{3}]y_{n}+[x_{n},x_{3}]y_{2})y_{1}^{i_{1}},x_{2},\ldots,x_{n})(x_{1}+[x_{n},x_{3}]y_{1}^{i_{1}}y_{n},x_{2},\ldots,x_{n})$
is tame. Since the second factor
$(x_{1}+[x_{n},x_{3}]y_{1}^{i_{1}}y_{n},x_{2},\ldots,x_{n})=(x_{1}+[x_{2},x_{3}]y_{1}^{i_{1}}y_{2},x_{2},\ldots,x_{n})^{(2n)}=\tau^{(2n)}$
of this product is also tame it follows that its first factor
$\sigma=(x_{1}+([x_{2},x_{3}]y_{n}+[x_{n},x_{3}]y_{2})y_{1}^{i_{1}},x_{2},\ldots,x_{n})$
is tame. By (5), we get
$3[x_{2},x_{3}]a=([x_{2},x_{3}]y_{n}+[x_{n},x_{3}]y_{2})y_{1}^{i_{1}}+([x_{2},x_{3}]y_{n}+[x_{2},x_{n}]y_{3})y_{1}^{i_{1}}.$
Consequently,
$C(3a)=(x_{1}+([x_{2},x_{3}]y_{n}+[x_{n},x_{3}]y_{2})y_{1}^{i_{1}},x_{2},\ldots,x_{n})$
$\times(x_{1}+([x_{2},x_{3}]y_{n}+[x_{2},x_{n}]y_{3})y_{1}^{i_{1}},x_{2},\ldots,x_{n})=\sigma(\sigma^{-1})^{(23)}$
is tame. If the characteristic of $K$ is not $3$, then its conjugation by $(x_{1},3x_{2},x_{3},\ldots,x_{n})$ gives $C(a)$.
If $n\geq 5$, then conjugating $C(a)$ by $(4n)$, we can make $i_{4}>0$. Case 1 gives that $C(a)$ is tame again.
Case 3. Let $i_{2}=i_{3}=\ldots=i_{n-1}=i_{n}=0$. In this case $a=y_{1}^{i_{1}}$ and $s=i_{1}\geq 2$ since $\mathrm{ldeg}(a)\geq 2$. Case 2 gives that
$\phi=(x_{1}+[x_{2},x_{3}]y_{1}^{s-1}y_{n},x_{2},\ldots,x_{n})$
is tame. Set $\beta=(x_{1},,\ldots,x_{n-1},x_{n}+x_{1})$. Direct calculations give that
$[\beta,\phi]=C(a)(x_{1},\ldots,x_{n-1},x_{n}-[x_{2},x_{3}]y_{1}^{s-1}(y_{1}+y_{n})$
$=C(a)(x_{1}-[x_{2},x_{3}]y_{n}^{s},\ldots,x_{n})^{(1n)}(x_{1}-[x_{2},x_{3}]y_{1}y_{n}^{s-1},x_{2},\ldots,x_{n})^{(1n)}.$
Since the second factor of this product is tame and the third factor is tame by Case 2 it follows that $C(a)$ is tame. $\Box$
This theorem contradicts the results of [39].
Corollary 1.
The group of all almost tame automorphisms $\mathrm{ATAut}(M_{n})$ is generated, modulo all linear automorphisms, by the quadratic automorphism (2) and the cubic automorphism (4).
Proof. The group of all tame automorphisms $\mathrm{TAut}(M_{n})$ is generated, modulo all linear automorphisms, by the quadratic automorphism (2) [35]. Obviuosly, the group of all almost tame automorphisms $\mathrm{ATAut}(M_{n})$ is generated, modulo all linear automorphisms, by Chein automorphisms $C(a)$. We have
$C(a)=C(a_{0})C(a_{1})\ldots C(a_{k}),$
where $a_{i}\in U_{i}$ for all $0\leq i\leq k$. We have $C(a_{0})$ is elementary since $a_{0}\in K$ and $C(a_{2}),\ldots,C(a_{k})$ are tame by Theorem 1.
Let
$a_{1}=\alpha_{1}y_{1}+\alpha_{2}y_{2}+\ldots+\alpha_{n}y_{n}.$
Then
$C(a_{1})=C(\alpha_{1}y_{1})C(\alpha_{2}y_{2})\ldots C(\alpha_{n}y_{n}).$
All $C(\alpha_{2}y_{2}),\ldots,C(\alpha_{n}y_{n})$ are again tame.
Consider $C(\alpha_{1}y_{1})$. Conjugating this by $(x_{1},\alpha_{1}^{-1}x_{2},x_{3},\ldots,x_{n})$ we can make $\alpha_{1}=1$ if $\alpha_{1}\neq 0$.
Consequently, the group of all almost tame automorphisms $\mathrm{ATAut}(M_{n})$ modulo $\mathrm{TAut}(M_{n})$ is generated by $C(y_{1})$. Note that $C(y_{1})$ is the cubic automorphism (4). Therefore the group of all almost tame automorphisms $\mathrm{ATAut}(M_{n})$ is generated, modulo all linear automorphisms, by the quadratic automorphism (2) and the cubic automorphism (4). $\Box$
@@ 4. Tame automorphisms
Every automorphism of $M_{n}$ of the form
$D(a)=(x_{1}+[x_{1},x_{2}]y_{1}a,x_{2}+[x_{1},x_{2}]y_{2}a,x_{3},\ldots,x_{n})$
is tame if $\mathrm{ldeg}(a)\geq 2$.
Proof. By Theorem 1, if $\mathrm{ldeg}(b)\geq 1$, then
$\phi=(x_{1}-[x_{3},x_{4}]y_{1}b,x_{2},\ldots,x_{n})(x_{1},x_{2}-[x_{3},x_{4}]y_{2}b,x_{3},\ldots,x_{n})$
$=(x_{1}-[x_{3},x_{4}]y_{1}b,x_{2}-[x_{3},x_{4}]y_{2}b,x_{3},\ldots,x_{n})$
is tame. Let
$\psi=(x_{1},x_{2},x_{3}+[x_{1},x_{2}],x_{4},\ldots,x_{n}).$
Then
$[\phi,\psi]=(x_{1}+[x_{1},x_{2}]y_{1}y_{4}b,x_{2}+[x_{1},x_{2}]y_{2}y_{4}b,x_{3},\ldots,x_{n})=D(y_{4}b)$
is tame.
Consequently, the statement of the lemma is true if $y_{4}$ divides $a$. Obviously, the same is true if $y_{i}$ divides $a$ for all $3\leq i\leq n$. Since
$D(a_{1}+a_{2})=D(a_{1})+D(a_{2}),\ \ a_{1},a_{2}\in U,$
we may assume that $a\in K[y_{1},y_{2}]$. Let $\alpha=(x_{1},x_{2},x_{3},x_{4}+x_{1},x_{5},\ldots,x_{n})$. If $b\in K[y_{1},y_{2}]$ then
$D(y_{4}b)^{\alpha}=(x_{1}+[x_{1},x_{2}]y_{1}(y_{4}+y_{1})b,x_{2}+[x_{1},x_{2}]y_{2}(y_{4}+y_{1})b,$
$x_{3},x_{4}-[x_{1},x_{2}]y_{2}(y_{4}+y_{1}),x_{5},\ldots,x_{n})$
$=D(y_{4}b)D(y_{1}b)(x_{1},x_{2},x_{3},x_{4}-[x_{1},x_{2}]y_{2}(y_{4}+y_{1}),x_{5},\ldots,x_{n}).$
Since $D(y_{4}b)$ is tame and
$(x_{1},x_{2},x_{3},x_{4}-[x_{1},x_{2}]y_{2}(y_{4}+y_{1}),x_{5},\ldots,x_{n})$
is tame by Theorem 1, this implies that $D(y_{1}b)$ is also tame. Therefore the statement of the lemma is true if $a\in K[y_{1},y_{2}]$ and $y_{1}$ divides $a$.
Changing $\alpha$ to $\beta=(x_{1},x_{2},x_{3},x_{4}+x_{2},x_{5},\ldots,x_{n})$ and repeating the same discussions as above, we can show that $D(y_{2}b)$ is also tame. $\Box$
Theorem 2.
Every exponential automorphism of $M_{n}$ of lower degree $\geq 5$ is tame.
Proof. Let
$E(m)=\exp(\mathrm{ad}(m))=(x_{1}+[m,x_{1}],x_{2}+[m,x_{2}],\ldots,x_{n}+[m,x_{n}])$
be an exponential automorphism with $\mathrm{ldeg}(m)\geq 4$. Since
$E(m_{1}+m_{2})=E(m_{1})+E(m_{2}),$
we may assume that $m=[x_{1},x_{2}]a$ with $\mathrm{ldeg}(a)\geq 2$. Then
$E(m)=E([x_{1},x_{2}]a)=D(a)\phi_{3}\ldots\phi_{n},$
where
$\phi_{i}=(x_{1},\ldots,x_{i-1},x_{i}+[x_{1},x_{2}]y_{i}a,x_{i+1},\ldots,x_{n}),\ 3\leq i\leq n.$
All $\phi_{i}$ are tame by Theorem 1 and $D(a)$ is tame by Lemma 4. Consequently, $E(m)$ is tame. $\Box$
This theorem contradicts the main result of [5].
5. Almost tame automorphisms
In this section we prove that a huge class of automorphisms of $M_{n}$ that moves only two variables $x_{1}$ and $x_{2}$ are almost tame. We always assume that $n\geq 4$ and $K$ is an arbitrary field. Since all Chein automorphisms are automatically included into $\mathrm{ATAut}(M_{n})$ by the definition, we don’t need any restrictions to characteristic of the field from the preceeding section.
First we give minor extensions of Lemma 4 and Theorem 2 for almost tame automorphisms.